Consider the following abbreviated financial statements for Cabo Wabo, Inc. CABO WABO, INC. Partial Balance Sheets as of December 31, 2015 and 2016 2015 2015 2016 Assets Current assets Net fxed aasets Liabilities and owners Equity $ 2,697$ 2,848 Current liabilites 12,411 2,969 Long-term debt S 1,135 S 1,679 6,687 7,846 CABO WABO, INC 2016 Income Slatement Sales Costs $39,950 20,026 3,395 Interest paid a. What is ownersequity for 2015 and 2016? (Do not round intermediate calculations and round your answers to the nearest whole number, e.g., 32.) Owners equity 2015 2016 b. What is the change in net working capital for 2016? (A negative answer should be indicated by a minus sign. Do not round intermediate calculations and round your answer to the nearest whole number, e.g, 32.) Change in net working capital c. In 2016, the company purchased $5,716 in new fixed assets. The tax rate is 40 percent. 1. How much in fixed assels did the company selr? (Do not round intermediate calculations and round your answer to the nearest whole number, e.g, 32.) Fixed assels sold 2. What is the cash flow from assets for the year? (Do not round intermediate calculations and round your answer to the nearest whole number, e.g., 32.) Cash flow from assets d. During 2016, the company raised $1,750 in new long-term debt. 1. What is the cash flow to creditors? (A negative answer should be indicated by a minus sign. Do not round intermediate calculations and round your answer to the nearest whole number, e.g., 32.) Cash flow to creditors 2. How much long-term debt must the company have paid off during the year? (Do not round intermediate calculations and round your answer to the nearest whole number, e.g., 32.) Debt retired
Answer #1

a) Owner's Equity in 2015 = Total Assets - total liabilities = (2697+12411)-(1135+6687) = 7268
Owner's Equity in 2016 = Total Assets - total liabilities = (2848+12969)-(1679+7846) = 6292

b) Change in NWC in 2016 = Working capital in 2016 - Working capital in 2015 = (2848-1679)-((2697-1135) = -393

c)

1) Fixed asset 2016 = Fixed assets 2015 - Depreciation + Asset bought - Asset Sold
12969 = 12411 - 3395 + 5716 - Assets Sold
Assets Sold = 1763

2) TAX = (Sales- COGS -Depreciation - Interest) * Tax rate = (39,959 - 20,026 -3395 - 636) *40% = 6357.20
Cash flow form Operations = EBIT -Taxes + Depreciation = (Sales - COGS- Depreciation) -Taxes + Depreciation = 13,566.80
  
Cash flow Assets = Cash Flow from Operations - Change in Fixed Assets - Change in NWC = 13566.80 - (12969-12411) -(-393) = 13401.80

d) Cash flow to creditors = Interest paid + New Debt = 636 + 1750 = 2386

Amount of Debt paid Off = Debt in 2015 - Debt in 2016 + New Debt = 6687 -7846 + 1750 = 591

Please Discuss in case of Doubt

Best of Luck. God Bless
Please Rate Well